# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v1

@@ II.2 Press-Schechter-type formalism
The estimation (6) neglect some important facts in the calculation of PBH abundance, which might change the result significantly.
@@ II.2.1 Compaction function
The criterion of PBH formation is not simply a threshold of the density contrast $\delta\rho/\rho$ Numerical study on the black hole collapse show that the criterion should be put on the averaged mass excess inside a sphere of a fixed areal radius, dubbed the compaction function Shibata and Sasaki (1999). In a reduced-circumference polar coordinates, we have  [^2] In this paper, following Escrivà et al. (2022b); Germani and Sheth (2023), we write the spatial metric as $\mathrm{d}\Sigma^{2}=e^{2\mathcal{R}(r)}\left(\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2}\right)=\frac{\mathrm{d}\bar{r}^{2}}{1-K(r)\bar{r}^{2}}+\bar{r}^{2}\mathrm{d}\Omega^{2}$ (7) where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}$ is the two-dimensional line element on a sphere. $\bar{r}$ is called reduced-circumference radius, which is related to $r$ by $\bar{r}=e^{\mathcal{R}(r)}r.$ (8)
$\mathscr{C}(t,\bar{r},\mathbf{x})\equiv\frac{2\delta M(r,t,\mathbf{x})}{r}=\frac{3}{\bar{r}}\left(aH\right)^{2}\int^{\bar{r}}_{0}\mathrm{d\varrho}~{}\varrho^{2}\delta(t,\varrho,\mathbf{x}).$ (9)
We can see that the compaction function is the averaged density contrast inside a sphere of radius $\bar{r}$, which is conserved on superhorizon scales. For the discussion of gauge issue, see Harada et al. (2015). It seems that, if we substitute the linear relation (3) into (9), we have
@@ II.2.2 Threshold
Around a density peak, the compaction function $\mathscr{C}(r)$ is a function of the radius $r$. Numerical calculations tell us that whether an overdensed region can collapse into a black hole depends on whether the maximum of the compaction function, $\mathscr{C}(r_{m})$, exceeds the critical value $\mathscr{C}_{\mathrm{th}}$, which varies from 0.4 to 0.67, depending on the density profile of the overdensed region. It was further found that only the width of the compaction function peak, parameterized by $q\equiv-\mathscr{C}^{\prime\prime}(r_{m})r_{m}^{2}/(4\mathscr{C}(r_{m}))$, is important Musco (2019). Large $q$ corresponds to a narrow $\mathscr{C}(r_{m})$, given by a top-hat density profile which drops rapidly around $r_{m}$. Strong pressure gradients at $r_{m}$ can resist the gravitational collapse, thus $\mathscr{C}_{\mathrm{th}}$ reaches its maximum $2/3$ in this limit. On the contrary, small $q$ corresponds to a broad $\mathscr{C}(r_{m})$, given by a density profile highly peaked at the center. As the pressure gradients can be neglected, it degenerate to the HYK limit $\mathscr{C}_{\mathrm{th}}\approx\delta_{\mathrm{th}}\approx 0.4$. In principle the density profiles and their statistical properties should be input if we want to calculate PBH abundance appropriately. Some hypothetical profiles are given by fitting the numerical results Musco (2019); Escrivà et al. (2020); Young (2019). In the peak theory, the profiles can be derived by the multipole moments of the power spectrum, provided the curvature perturbation is Gaussian Yoo et al. (2018); Atal et al. (2019, 2020a); Yoo et al. (2021); Kitajima et al. (2021).
@@
We see that depending on whether $\mathscr{C}_{\ell}<4/3$ or $\mathscr{C}_{\ell}>4/3$, the maximum $\mathscr{C}(r_{m})$ corresponds to the maximum or minimum of $\mathscr{C}_{\ell}$. Only the former case, dubbed type I PBH Kopp et al. (2011), is consistent with the linear approximation. Type II PBHs with $\mathscr{C}_{\ell}>4/3$ are rarer, which we do not consider in this paper. Given the profile-dependent threshold on $\mathscr{C}$, the corresponding threshold on the linear compaction function $\mathscr{C}_{\ell}$ can be solved by (12)
@@
As two examples, the threshold of a top-hat profile $\mathscr{C}_{\mathrm{th}}\approx 1/2$ gives $\mathscr{C}_{\ell,\mathrm{th}}\approx 2/3$. In the peak theory, a monochromatic power spectrum gives $\mathscr{C}_{\mathrm{th}}\approx 0.587$, thus $\mathscr{C}_{\ell,\mathrm{th}}\approx 0.872$ Kitajima et al. (2021).
When the power spectrum of $\mathcal{R}$ is broad, for every scale we study, we should add a window function to smooth out the unwanted short-wavelength fluctuations. This is highly non-trivial and there will be uncertainties from the choice of the window function, which will not be discussed in this paper.
@@ II.2.3 Primordial non-Gaussianity
We already know from Section II.2.1 that the compaction function $\mathscr{C}$ is a quadratic function of $\mathscr{C}_{\ell}$. In the presence of primordial non-Gaussianity in $\mathcal{R}$, even $\mathscr{C}_{\ell}$ is non-Gaussian. As we will see, in typical cases, $\mathcal{R}$ is a function of a Gaussian-distributed variable, $\mathcal{R}_{\mathrm{g}}$, which obeys the Gaussian PDF
@@
$\sigma_{\mathcal{R}}^{2}=\int\frac{\mathrm{d}k}{k}\tilde{W}^{2}(k,r)\mathcal{P}_{\mathcal{R}_{g}}(k).$ (16)
For simplicity, let’s assume that the non-Gaussianity is local-type, which gives
@@ II.3 PBH abundance with primordial non-Gaussianity
where PNG stands for primordial non-Gaussianity of the curvature perturbation generated during inflation, which will be discussed in details in the next section. Calculations following the routines in (20) are called extended Press-Schechter or Press-Schechter-type formalism, which is in principle different from another frequently used method, the peak theory Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022), by two main points. In extended Press-Schechter fomalism, the PBH abundance is calculated by the integrating the PDF of $\mathscr{C}_{\ell}$ from the threshold given by $\mathscr{C}_{\ell}\geq\mathscr{C}_{\ell,\mathrm{th}}$, while in the peak theory the PBH abundance is given by counting the number of peaks $n_{\mathrm{pk}}$ where the averaged compaction function in a sphere of radius $r_{m}$, $\langle\mathscr{C}\rangle_{r\leq r_{m}}$, exceeds a universal threshold 2/5 Escrivà et al. (2020).
@@ III Primordial non-Gaussianities of the curvature perturbation
In slow-roll inflation, the nonlinear parameter $f_{\mathrm{NL}}$ is proportional to the spectral tilt, i.e. the consistency relation $f_{\mathrm{NL}}=(5/12)(n_{s}-1)$ Maldacena (2003); Creminelli and Zaldarriaga (2004), which is suppressed by slow-roll parameters as $n_{s}-1=-2\epsilon_{H}-\eta_{H}$. On scales larger than 1 Mpc, the observational constraint is $f_{\mathrm{NL}}=-0.9\pm 5.1$ Akrami et al. (2020b); Abbott et al. (2022).
As is commented above, the PBH abundance depends sensitively on the non-Gaussianities, thus even a small non-Gaussianity can change the PBH abundance significantly. Furthermore, as the power spectrum is enhanced rapidly (usually as power-law) before the peak scale, an intuitive guess based on the consistency relation $f_{\mathrm{NL}}\sim n_{s}-1$ is that non-Gaussianity will also be enhanced to $\mathcal{O}(1)$. For instance, in ultra-slow-roll inflation, $\epsilon_{H}\lll 1$ and $\eta_{H}=-6$. If we only focus on the slow-roll attractor mode, we have
$f_{\mathrm{NL}}=\frac{5}{12}(n_{s}-1)=-\frac{5}{12}(2\epsilon_{H}+\eta_{H})\approx-\frac{5}{12}\times(-6)=\frac{5}{2}.$ (22)
Astonishingly, it gives the correct value, although it is in contradiction with the consistency relation of the total power spectrum $\mathcal{P_{R}}(k)\propto k^{4}$, of which the slope originates from the next-to-leading order of gradient expansion Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022).
In principle, when $f_{\mathrm{NL}}\gtrsim 1$, the higher order terms includes an infinite number of parameters, which makes it impossible to have a model-independent analysis  [^3] In some specific models, it is possible to truncate the perturbative series at quadratic level, no matter how large $f_{\mathrm{NL}}$ is. See Section III.4 for details.. Fortunately, the non-Gaussian curvature perturbation can be calculated by $\delta N$ formalism, based on the separate universe approach Sasaki and Stewart (1996); Wands et al. (2000); Lyth et al. (2005). During inflation, the gradient terms decay rapidly on superhorizon scales, which guarantees that in a region slightly larger than the Hubble patch, the inflaton field takes a homogeneous value, which varies from region to region. The initial condition for $\varphi$ in the numerous separate universes obeys a Gaussian distribution, which evolves independently on superhorizon scales. $\delta N$ formalism tells us that the final curvature perturbation in such a patch is the difference of total local expansion, i.e. the $e$-folding number $N$, between this patch to the fiducial $\langle N\rangle$, evaluated from an initial spatially-flat slice to the final comoving slice:
@@
This formula does not require a decomposition of a background and a perturbation, thus can be used for large perturbations. We will use it to study the fully nonlinear curvature perturbation in the following sections. In Vennin and Wands (2024), (23) will be used to study the quantum diffusion.
@@ III.1 Ultra-slow-roll inflation on a flat potential
When $f_{\mathrm{NL}}\gtrsim 1$, we should take into account all the higher order terms, which is only possible for some specific models. The simplest example is the so-called ultra-slow-roll inflation. Consider a plateau of potential $V(\varphi)=V_{0}$ for $\varphi_{i}<\varphi<\varphi_{t}$. On the plateau, the equation of motion of $\varphi$ is $\ddot{\varphi}+3H\dot{\varphi}=0$, where $H\approx\sqrt{V_{0}}/(\sqrt{3}M_{\mathrm{Pl}})$ is nearly a constant on the plateau. Conventionally, we define the $e$-folding number $N$, counted backward in time from $N(\varphi_{t})=0$, such that $\mathrm{d}N=-H\mathrm{d}t$. The equation of motion becomes
@@
Two general solutions to (24) are $\varphi=\mathrm{constant}$ and $\varphi\propto e^{3N}$. The former static solution is the attractor, but it should not dominate at the classical level, otherwise the inflaton gets stuck on the plateau, which can only end with the help of quantum diffusion. The non-attractor decelerating solution $\varphi\propto e^{3N}$ is necessary as it helps the inflaton to classically shoot out of the potential plateau with large enough initial velocity. We will see that the excitation of such non-attractor solutions is the origin of the large non-Gaussianity. Together with the boundary condition at the endpoint $\varphi_{t}$ of the plateau, the special solution is
@@
which is valid for all the trajectories at any $e$-folding number. For a perturbed universe, (26) gives $\tilde{\pi}(N)+3\tilde{\varphi}(N)=\tilde{\pi}_{t}+3\varphi_{t}$, which in turn gives
$\delta\pi(N)+3\delta\varphi(N)=\delta\pi_{t}.$ (27)
We defined $\delta\varphi=\tilde{\varphi}-\varphi$, $\delta\pi\equiv\tilde{\pi}-\pi$ and $\delta\pi_{t}\equiv\tilde{\pi}_{t}-\pi_{t}$, which are Gaussian variables as $\varphi$ is a massless field. The physical meaning of (26) and (27) can be well understood in the phase portrait Fig. 1.
Figure 1: The schematic phase portrait of the ultra-slow-roll inflation on a flat plateau. The blue and green lines with arrows are the inflaton trajectory in the background universe and a perturbed universe, respectively. The red lines are equal-$N$ lines, determined by the conservation law (26), while ($\tilde{\pi}$, $\tilde{\varphi}$) at the perturbed trajectory should have the same $N$ as ($\pi$, $\varphi$), i.e. the perturbation is along an equal-$N$ line. Such a perturbed point ($\tilde{\pi}$, $\tilde{\varphi}$) satisfies (27), as is clearly shown geometrically. For a detailed discussion of such a phase portrait of ultra-slow-roll inflation, see Passaglia et al. (2019).
@@
In the separate universe approach, another remote and independent “perturbed universe” has expanded another $e$-folding number of $\tilde{N}=(1/3)\ln\tilde{\pi}/\tilde{\pi}_{t}$, with $\tilde{\pi}=\pi+\delta\pi$, and $\tilde{\pi}_{t}=\pi_{t}+\delta\pi_{t}$. In the $\delta N$ formalism, the late-time curvature perturbation in this perturbed universe is the difference between these $e$-folding numbers
$\delta N=\tilde{N}-N=\frac{1}{3}\ln\frac{\tilde{\pi}}{\tilde{\pi}_{t}}\frac{\pi_{t}}{\pi}=\frac{1}{3}\ln\left(1-\frac{\delta\pi}{\pi}\right)-\frac{1}{3}\ln\left(1-\frac{\delta\pi_{t}}{\pi_{t}}\right)$ (29)
Note that from from (27) we have $\delta\pi\sim\delta\pi_{t}$, so unless $N\approx 0$, the first term in (29) is always neglegible compared to the second term as $\pi=\pi_{t}e^{3N}\gg\pi_{t}$. If the later evolution do not contribute significantly to $\delta N$, [^4] This requirement is not trivial. It is equivalent to require that the inflaton accelerates to the later slow-roll attractor after it classically shoots out of the flat pleateau. On the other hand, if $\varphi$ decelerates, the $\delta N$ contributed from the later stage is dominant, which is usually Gaussian and thus the entire curvature perturbation is nearly Gaussian. See below. the final curvature perturbation is

### v2

@@ II.2 Press-Schechter-type formalism
The estimation (6) neglects some important facts in the calculation of PBH abundance, which might change the result significantly.
@@ II.2.1 Compaction function
The criterion of PBH formation is not simply a threshold of the density contrast $\delta\rho/\rho$. Numerical studies on the black hole collapse show that the criterion should be put on the averaged mass excess inside a sphere of a fixed areal radius, dubbed the compaction function Shibata and Sasaki (1999). In a reduced-circumference polar coordinates, we have  [^3] In this paper, following Refs. Germani and Sheth (2020); Escrivà et al. (2022b); Germani and Sheth (2023), we write the spatial metric as $\mathrm{d}\Sigma^{2}=e^{2\mathcal{R}(r)}\left(\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2}\right)=\frac{\mathrm{d}\bar{r}^{2}}{1-K(\bar{r})\bar{r}^{2}}+\bar{r}^{2}\mathrm{d}\Omega^{2}$ (7) where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}$ is the two-dimensional line element on a sphere. $\bar{r}$ is called reduced-circumference radius, which is related to $r$ by the local expansion: $\bar{r}=e^{\mathcal{R}(r)}r.$ (8)
$\mathscr{C}(t,\bar{r},\mathbf{x})\equiv\frac{2\delta M(t,\bar{r},\mathbf{x})}{\bar{r}}=\frac{3}{\bar{r}}\left(aH\right)^{2}\int^{\bar{r}}_{0}\mathrm{d\varrho}~{}\varrho^{2}\delta(t,\varrho,\mathbf{x}),$ (9)
where we set $G=1$. We can see that the compaction function is the averaged density contrast inside a sphere of radius $\bar{r}$, which is conserved on superhorizon scales. For the discussion of gauge issue, see Harada et al. (2015). It seems that, if we substitute the linear relation (3) into (9), we have
@@ II.2.2 Threshold
Around a density peak, the compaction function $\mathscr{C}(\bar{r})$ is a function of the radius $\bar{r}$. Numerical calculations tell us that whether an overdensed region can collapse into a black hole depends on whether the maximum of the compaction function, $\mathscr{C}(\bar{r}_{m})$, exceeds the critical value $\mathscr{C}_{\mathrm{th}}$, which varies from 0.41 to 0.67, depending on the density profile of the overdensed region Musco (2019). It was further found that only the width of the compaction function peak, parameterized by $q\equiv-\mathscr{C}^{\prime\prime}(\bar{r}_{m})\bar{r}_{m}^{2}/(4\mathscr{C}(\bar{r}_{m}))$, is important Escrivà et al. (2020). Large $q$ corresponds to a narrow $\mathscr{C}(\bar{r}_{m})$, given by a top-hat density profile which drops rapidly around $\bar{r}_{m}$. Strong pressure gradients at $\bar{r}_{m}$ can resist the gravitational collapse, thus $\mathscr{C}_{\mathrm{th}}$ reaches its maximum $2/3$ in this limit. On the contrary, small $q$ corresponds to a broad $\mathscr{C}(\bar{r}_{m})$, given by a density profile highly peaked at the center. As the pressure gradients can be neglected, it reduces to the HYK limit $\mathscr{C}_{\mathrm{th}}\approx\delta_{\mathrm{th}}\approx 0.4$. In principle the density profiles and their statistical properties should be input if we want to calculate PBH abundance appropriately. Some hypothetical profiles are given by fitting the numerical results Musco (2019); Escrivà et al. (2020); Young (2019). In the theory of peaks, the profiles can be derived by the multipole moments of the power spectrum, provided the curvature perturbation is Gaussian Yoo et al. (2018); Atal et al. (2019, 2020a); Yoo et al. (2021); Kitajima et al. (2021).
@@
We see that depending on whether $\mathscr{C}_{\ell}<4/3$ or $\mathscr{C}_{\ell}>4/3$, the maximum $\mathscr{C}(r_{m})$ corresponds to the maximum or minimum of $\mathscr{C}_{\ell}$. Only the former case, dubbed type I PBH Kopp et al. (2011), is consistent with the linear approximation. Type II PBHs with $\mathscr{C}_{\ell}>4/3$ are usually rarer, which we do not consider in this paper. Given the profile-dependent threshold on $\mathscr{C}$, the corresponding threshold on the linear compaction function $\mathscr{C}_{\ell}$ can be solved by (12)
@@
As two examples, the threshold of a Maxican-hat profile considered in Chapter 7 Young (2024) is $\mathscr{C}_{\mathrm{th}}\approx 1/2$, which gives $\mathscr{C}_{\ell,\mathrm{th}}\approx 2/3$. In the theory of peaks, a monochromatic power spectrum gives $\mathscr{C}_{\mathrm{th}}\approx 0.587$, thus $\mathscr{C}_{\ell,\mathrm{th}}\approx 0.872$ Kitajima et al. (2021), which we will use later.
When the power spectrum of $\mathcal{R}$ is broad, for every scale we study, we should add a window function to smooth out the unwanted short-wavelength fluctuations. This is highly non-trivial and there are uncertainties from the choice of the window function, which will not be discussed in this paper.
@@ II.2.3 Primordial non-Gaussianity
We already know from Section II.2.1 that the compaction function $\mathscr{C}$ is a quadratic function of $\mathscr{C}_{\ell}$. In the presence of primordial non-Gaussianity in $\mathcal{R}$, even $\mathscr{C}_{\ell}$ is non-Gaussian. As we will see, for local non-Gaussianities, $\mathcal{R}$ is a function of a Gaussian-distributed variable, $\mathcal{R}_{\mathrm{g}}$, which obeys the Gaussian PDF
@@
$\sigma_{\mathcal{R}}^{2}=\int\frac{\mathrm{d}k}{k}\tilde{W}^{2}(k,r)\mathcal{P}_{\mathcal{R}\mathrm{g}}(k).$ (16)
Local-type non-Gaussianity gives
@@ II.3 PBH abundance with primordial non-Gaussianity
where PNG stands for primordial non-Gaussianity of the curvature perturbation generated during inflation, which will be discussed in details in the next section. Calculations following the routines in (20) are called extended Press-Schechter or Press-Schechter-type formalism, which is in principle different from another frequently used method, the theory of peaks Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022), by two main points. In Press-Schechter-type formalism, the PBH abundance is calculated by integrating the PDF of $\mathscr{C}_{\ell}$ from the threshold given by $\mathscr{C}_{\ell}\geq\mathscr{C}_{\ell,\mathrm{th}}$, while in the theory of peaks the PBH abundance is given by counting the number of peaks $n_{\mathrm{pk}}$ where the averaged compaction function in a sphere of radius $r_{m}$, $\langle\mathscr{C}\rangle_{r\leq r_{m}}$, exceeds a universal threshold 2/5 Escrivà et al. (2020).
@@ III Primordial non-Gaussianities of the curvature perturbation
In slow-roll inflation, the nonlinear parameter $f_{\mathrm{NL}}$ is related to the spectral tilt by the consistency relation $f_{\mathrm{NL}}=(5/12)(1-n_{s})$ Maldacena (2003); Creminelli and Zaldarriaga (2004), which is suppressed by slow-roll parameters as $n_{s}-1=-2\epsilon_{H}-\eta_{H}$, with $\epsilon_{H}\equiv-\dot{H}/H^{2}$ and $\eta_{H}\equiv\dot{\epsilon}_{H}/(H\epsilon_{H})$. On scales larger than 1 Mpc, the observational constraint is $f_{\mathrm{NL}}=-0.9\pm 5.1$ Akrami et al. (2020b); Abbott et al. (2022).
As is commented above, the PBH abundance depends sensitively on the non-Gaussianities, thus even a small non-Gaussianity can change the PBH abundance significantly. Furthermore, as the power spectrum is enhanced rapidly (usually as power-law, $k^{\mathcal{O}(1)}$) before the peak scale, an intuitive guess based on the consistency relation $f_{\mathrm{NL}}\sim 1-n_{s}$ is that non-Gaussianity will also be enhanced to $\mathcal{O}(1)$. For instance, in ultra-slow-roll inflation, $\epsilon_{H}\lll 1$ and $\eta_{H}=-6$. If we only focus on the slow-roll attractor mode, we have
$f_{\mathrm{NL}}=\frac{5}{12}(1-n_{s})=\frac{5}{12}(2\epsilon_{H}+\eta_{H})\approx\frac{5}{12}\times(-6)=-\frac{5}{2}.$ (22)
Accidentally, this rough estimate gives the correct absolute value with an opposite sign Namjoo et al. (2013). Anyway this indicates that the non-Gaussianity may reach $\mathcal{O}(1)$ when the power spectrum increases rapidly.
In principle, when $|f_{\mathrm{NL}}|\gtrsim 1$, the higher order terms include an infinite number of parameters, i.e. $g_{\mathrm{NL}}$, $h_{\mathrm{NL}}$, $i_{\mathrm{NL}}$, etc., which makes it impossible to have a model-independent analysis. Fortunately, in some typical models, the non-Gaussian curvature perturbation can be calculated neatly by $\delta N$ formalism, based on the separate universe approach Sasaki and Stewart (1996); Wands et al. (2000); Lyth et al. (2005). During inflation, the gradient terms decay rapidly on superhorizon scales, which guarantees that in a region slightly larger than the Hubble patch, the inflaton field takes a homogeneous value, which varies from region to region. The initial condition for $\varphi$ in the numerous separate universes obeys a Gaussian distribution, which evolves independently on superhorizon scales. $\delta N$ formalism tells us that the final curvature perturbation in such a patch is the difference of total local expansion, i.e. the $e$-folding number $N$, between this patch to the fiducial $\langle N\rangle$, evaluated from an initial spatially-flat slice to the final comoving slice:
@@
This formula does not require a decomposition of a background and a perturbation, thus can be used for large perturbations. We will use it to study the fully nonlinear curvature perturbation in the following sections. In Chapter 9 Vennin and Wands (2024), (23) will be used to study the quantum diffusion.
@@ III.1 Ultra-slow-roll inflation on a flat potential
When $|f_{\mathrm{NL}}|\gtrsim 1$, we should take into account all the higher order terms, which is only possible for some specific models. The simplest example is the so-called ultra-slow-roll inflation. Consider a plateau of potential $V(\varphi)=V_{0}$ for $\varphi_{i}<\varphi<\varphi_{t}$. On the plateau, Inflaton with an initial velocity moves toward the positive-$\varphi$ direction. The equation of motion of $\varphi$ is $\ddot{\varphi}+3H\dot{\varphi}=0$, where $H\approx\sqrt{V_{0}}/(\sqrt{3}M_{\mathrm{Pl}})$ is nearly a constant on the plateau ($M_{\mathrm{Pl}}=1/\sqrt{8\pi G}$). Conventionally, we define the $e$-folding number $N$, counted backward in time from $N(\varphi_{t})=0$, such that $\mathrm{d}N=-H\mathrm{d}t$. The equation of motion becomes
@@
Two general solutions to (24) are $\varphi=\mathrm{constant}$ and $\varphi\propto e^{3N}$. The former static solution is the attractor, but it should not dominate at the classical level, otherwise the inflaton gets stuck on the plateau, which can only end inflation with the help of quantum diffusion. The non-attractor decelerating solution $\varphi\propto e^{3N}$ is necessary as it helps the inflaton to classically shoot out of the potential plateau with large enough initial velocity. We will see that the excitation of such non-attractor solutions is the origin of the large non-Gaussianity. Together with the boundary condition at the endpoint $\varphi_{t}$ of the plateau, the special solution is
@@
which is valid for all the trajectories at any $e$-folding number. For a perturbed universe, (26) gives $\tilde{\pi}(\tilde{N})+3\tilde{\varphi}(\tilde{N})=\tilde{\pi}_{t}+3\varphi_{t}$, which in turn gives
$\delta\pi+3\delta\varphi=\delta\pi_{t}.$ (27)
We defined $\delta\varphi\equiv\tilde{\varphi}-\varphi$, $\delta\pi\equiv\tilde{\pi}-\pi$ and $\delta\pi_{t}\equiv\tilde{\pi}_{t}-\pi_{t}$, which are Gaussian variables as $\varphi$ is a massless field. The physical meaning of (26) and (27) can be well understood in the schematic phase portrait Fig. 1.
Figure 1: The schematic phase portrait of the ultra-slow-roll inflation on a flat plateau. The blue and green parallel lines with arrows are the inflaton trajectories in the background universe and in a perturbed Hubble patch, respectively. The red lines are equal-$N$ lines, determined by the conservation law (26), while ($\tilde{\pi}$, $\tilde{\varphi}$) at the perturbed trajectory corresponds to an $e$-folding number $\tilde{N}=N+\delta N$. Such a perturbed point ($\tilde{\pi}$, $\tilde{\varphi}$) satisfies (27), as is clearly shown geometrically. For more discussion on such a phase portrait, see Passaglia et al. (2019).
@@
In the separate universe approach, another remote and independent “perturbed universe” happens to have another $e$-folding number of $\tilde{N}=(1/3)\ln\tilde{\pi}/\tilde{\pi}_{t}$, with $\tilde{\pi}=\pi+\delta\pi$, and $\tilde{\pi}_{t}=\pi_{t}+\delta\pi_{t}$. In the $\delta N$ formalism, the late-time curvature perturbation in this perturbed universe is the difference between these $e$-folding numbers
$\delta N=\tilde{N}-N=\frac{1}{3}\ln\left(\frac{\tilde{\pi}}{\tilde{\pi}_{t}}\frac{\pi_{t}}{\pi}\right)=\frac{1}{3}\ln\left(1-\frac{\delta\pi}{\pi}\right)-\frac{1}{3}\ln\left(1-\frac{\delta\pi_{t}}{\pi_{t}}\right)$ (29)
Note that from from (27) we have $\delta\pi\sim\delta\pi_{t}$, so unless $N\approx 0$, the first term in (29) is always neglegible compared to the second term as $\pi=\pi_{t}e^{3N}\gg\pi_{t}$. If the later evolution do not contribute significantly to $\delta N$, [^4] This requirement is not trivial. It is equivalent to require that the inflaton accelerates to the later slow-roll attractor after it classically shoots out of the flat pleateau. On the other hand, if $\varphi$ decelerates, the $\delta N$ contributed from the later stage is dominant, which is usually Gaussian and thus the entire curvature perturbation is nearly Gaussian. See Section III.3 for a detailed discussion. the final curvature perturbation is